OpenStudy (anonymous):

find the root of the polynomial: a^2/16 + 3a+36

OpenStudy (anonymous):

a = -24

OpenStudy (anonymous):

-24 is the root??

OpenStudy (anonymous):

Yeah...

OpenStudy (anonymous):

\[\frac{ a^2{} }{ 16 } + 3a+36\]

OpenStudy (anonymous):

i don't get it,,,

OpenStudy (anonymous):

= 0, right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it is just find the root of the polynomial

OpenStudy (anonymous):

Then it must be equal to 0. a = -24 is the answer.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

-24^2 = 576 So (576/16)+3(-24)+36 = 0